# The $\theta$-vacuum from functional renormalisation

### v1

@@ D.1 Flow with topological freezing
In this Appendix we briefly review the fRG approach used in Fukushima:2022zor . There, the flow equation for the effective action in 87 was derived analogously to Appendix B. However, instead of the complex field basis a Cartesian basis was used. The field $\varphi$ was disentangled in its real and imaginary part, $\varphi=x+iy$. Using the Litim cutoff function 40, the flow equation for the effective potential takes the form
@@
Now we use the numerical solution of 88 with the initial potential 2 for studying the $\theta$-dependence of the ground-state energy. For a benchmark coupling value of $g=30$, the flow equation yields results for the ground-state energy that closely match those obtained from the Schrödinger equation within the regime $\theta\lesssim 2\pi$, where $n=0$ corresponds to the ground state. However, as $\theta$ increases beyond $2\pi$, the solution to the flow equation breaks down due to the emergence of a pole in the flow, leading to a termination of the flow of the ground-state energy around $\theta\sim 2\pi$. This is also seeing in Figure 4 in Fukushima:2022zor or the left-hand side panel in Figure 12.
@@ D.2 Flow with topological freezing and complex frequencies
We have also investigated a variant of the flow with topological freezing, where we exploited that fact, that the topological term can be absorbed in an imaginary shift of the frequency. This circumvents the pole observed in Section D.1 and potentially accommodates topological effects analogously to resurgence.
For this purpose we rewrite the kinetic and topological terms in the initial action 1 as
@@
where we defined the complex frequency
@@
This is analogous to the system with a chemical potential if $\mu\equiv\tilde{\theta}$. The potential receives an additional contribution proportional to $\theta^{2}$ from the second term in 89 and then the initial potential is given by
@@
In such a setup, the system may be interpreted as a system with the polynomial potential 91 at finite density. Here, we may make an ansatz for the effective action as
@@
In this case, the frequency integral of the action has been extended into the complex plane due to the complex momentum 90. This is similar to theories with a chemical potential. There, the Silver Blaze property requires the use of regulators that do not destroy the underlying complex structure. We have refrained from such an investigation here as it goes beyond the scope of the present work.
[…]
@@
$R_{k}(p_{\theta})=\operatorname{Re}(k^{2}-p_{\theta}^{2})\Theta\bigl{(}\operatorname{Re}(k^{2}-p_{\theta}^{2})\bigr{)}=\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,\Theta\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,.$ (93)
For such a regulator, we compute
$\partial_{t}R_{k}(p_{\theta})$ $=2k^{2}\,\Theta\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,,$ (94)
@@
$p_{\theta}^{2}+R_{k}(p_{\theta})$ $=p^{2}-\tilde{\theta}^{2}+2i\tilde{\theta}p+(k^{2}+\tilde{\theta}^{2}-p^{2})\Theta\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)$
$=2i\tilde{\theta}p+k^{2}\,\Theta\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)+\left(p^{2}-\tilde{\theta}^{2}\right)\,\Theta\left(p^{2}-k^{2}-\tilde{\theta}^{2}\right)\,.$ (95)
The last line in 95 does not contribute to the flow of the effective potential as it lies outside the range of the momentum integral with $p^{2}\leq k^{2}+\tilde{\theta}^{2}$ following from 94. The flow equation is given by
@@
in which the integral variable range reads $\Omega=\left[-\sqrt{k^{2}+\tilde{\theta}^{2}},\sqrt{k^{2}+\tilde{\theta}^{2}}\right]$. Performing the momentum integral and taking the real part, we obtain
@@
where $c^{(1)}_{k}$ and $c^{(2)}_{k}$ are defined in 85.
Now we solve the flow equation for $V_{k}(r^{2})$ with the UV boundary condition 91 at $k=\Lambda$. In the right-hand side panel of Figure 12, we show the result of the vacuum energy in terms of the $\theta$-parameter by setting $m=1$. As the same as the real momentum case, the energy level from the flow equation 97 breaks down for $\theta\gtrsim 2\pi$. This happens because at an intermediate energy scale we have $k^{2}+2V^{\prime}=0$ for which the right-hand side of 97 diverges.
We close this discussion with the remark that this setup deserves further analysis using regulators that keep the complex structure intact. We shall consider this elsewhere.
Appendix E RG evolution of effective potential
[…]
@@ Appendix F Flow of the cusp position
In this Appendix we derive relations for the flowing cups positions. The derivation is reminiscent of that of the flow of the flowing solution of the equations of motion. We keep this derivations minimal as we do not use these relations in the present work. Still we have checked their applicability numerically.
@@
The numerator is simply the third derivative of the flow, evaluated at $\varphi_{k}^{\textrm{cusp}}$ with 98. All terms except one drop out because they are proportional to powers of $V_{k}^{(3)}(\varphi_{k}^{\textrm{cusp}})$.
@@
We emphasise that 100 holds true for any field value but for all but $\varphi^{\textrm{cusp}}$ both side in 101 vanish identically. In turn, for $\varphi_{k}^{\textrm{cusp}}$ they both are non-vanishing and we arrive at
@@
One may be tempted to use 102 in the finite-volume case where we have have $V^{+}_{k}=V^{-}_{k}$, but this leads us to 101 with vanishing left and right hand side, and 102 does not follow. Indeed, ignoring this fact leads us from 102 to
@@
Equation 103 vanishes on the cusps in clear contradiction to the observed flow of $\varphi_{k}^{\textrm{cusp}}$ either given by 99 or being extracted from the results. The vanishing of the right hand side of 103 originates in the fact that the flow of $V^{(1)}(\varphi_{k}^{\textrm{cusp}})$ is proportional to $V^{(3)}(\varphi_{k}^{\textrm{cusp}})=0$ and 103 vanishes. This concludes our discussion of the flow of $\varphi_{k}^{\textrm{cusp}}$.

### v2

@@ D.1 Flow with topological freezing
In this Appendix we briefly review the fRG approach used in Fukushima:2022zor . There, the flow equation for the effective action in 88 was derived analogously to Appendix B. However, instead of the complex field basis a Cartesian basis was used. The field $\varphi$ was disentangled in its real and imaginary part, $\varphi=x+iy$. Using the Litim cutoff function 40, the flow equation for the effective potential takes the form
@@
For the numerical solution of 89 it is more convenient to consider the flow of the derivative of the potential $V_{k}^{\prime}(r^{2})\equiv\partial_{r^{2}}V_{k}(r^{2})$ rather than the potential $V_{k}(r^{2})$ itself: The $r^{2}$-derivative leads to higher order poles in the flow equation for $\Gamma^{(2)}+R_{k}\to 0$. These poles are repulsive and encode the convexity-restoring property of the Wetterich equation, see Litim:2006nn . While soft or even integrable singularities in the flow equation still accommodate convexity, the numerical access is far more challenging.
Consequently we consider
$\partial_{t}V_{k}^{\prime}(r^{2})=\frac{\partial}{\partial r^{2}}\Big[\partial_{t}V_{k}(r^{2})\Big]\,.$ (90)
This allows us to compute $V^{\prime}_{k=0}$ and the effective potential is obtained from an $r^{2}$-integration,
$V_{k}(r^{2})=\int_{r_{0}^{2}}^{r^{2}}\text{d}(r_{1}^{2})\ V_{k}^{\prime}(r_{1}^{2})+\textrm{const.}\,$, (91)
where the constant includes a potential shift such as used in Sections IV.2 and IV.1.
In the left-hand side panel of Figure 13, we show the result following the above strategy in the red line with markers, with $r_{0}^{2}=1$ as the boundary field value. Evidently, the breakdown of the numerical solution of the flow at $\theta=2\pi$, reported on in Fukushima:2022zor , is absent. Indeed, this breakdown can be traced back to the soft singularity structure of the flow $\partial_{t}V$ which is resolved by considering $\partial_{t}V^{\prime}$.
The result for $V(r)$ does not show the level crossing in line with the arguments in Section III.2: the flows do not accommodate topological configurations and hence the level crossing is absent. However the derivation still includes the topological term. It is suggestive that this term deforms the correct flow which leads to a deviation of the exact results with increasing $\theta$.
@@ D.2 Flow with topological freezing and complex frequencies
We have also investigated a variant of the flow with topological freezing: in this setup we exploit, that the topological term can be absorbed in an imaginary shift of the frequency. This potentially accommodates topological effects analogously to resurgence. For this purpose we rewrite the kinetic and topological terms in the initial action 1 as
@@
with the complex frequency
@@
This is analogous to the system with a chemical potential if $\mu\equiv\tilde{\theta}$. The potential receives an additional contribution proportional to $\theta^{2}$ from the second term in 92 and then the initial potential is given by
@@
In such a setup, the system may be interpreted as a system with the polynomial potential 94 at finite density. Here, we may make an ansatz for the effective action as
@@
In this case, the frequency integral of the action has been extended into the complex plane due to the complex momentum 93. This is similar to theories with a chemical potential. There, the Silver Blaze property requires the use of regulators that do not destroy the underlying complex structure.
We have refrained from such an investigation here as it goes beyond the scope of the present work.
Figure 13: Energy eigenvalues for $n=0,\,\cdots,\,3$. The dashed-circle line exhibits the ground-state energy in terms of the $\theta$-parameter from 89 (left) and 101 (right). The dashed-colored lines show the solutions to the Schrödinger equation 13, and the solid-black line denotes the result obtained from the flow equation with the Litim-type regulator.
@@
$R_{k}(p_{\theta})=\operatorname{Re}(k^{2}-p_{\theta}^{2})\Theta\bigl(\operatorname{Re}(k^{2}-p_{\theta}^{2})\bigr)=\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,\tilde{\Theta}\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,$, (96)
where
$\tilde{\Theta}\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)=\Theta\bigl(\operatorname{Re}(k^{2}-p_{\theta}^{2})\bigr)\,.$ (97)
Equation 97 implies
$\partial_{t}R_{k}(p_{\theta})$ $=2k^{2}\,\tilde{\Theta}\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)\,$, (98)
@@
$p_{\theta}^{2}+R_{k}(p_{\theta})$ $=p^{2}-\tilde{\theta}^{2}+2i\tilde{\theta}p+(k^{2}+\tilde{\theta}^{2}-p^{2})\tilde{\Theta}\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)$
$=2i\tilde{\theta}p+k^{2}\,\tilde{\Theta}\left(k^{2}+\tilde{\theta}^{2}-p^{2}\right)+\left(p^{2}-\tilde{\theta}^{2}\right)\,\tilde{\Theta}\left(p^{2}-k^{2}-\tilde{\theta}^{2}\right)\,.$ (99)
The last term of the second line in 99 does not contribute to the flow of the effective potential as it lies outside the range of the momentum integral with $p^{2}\leq k^{2}+\tilde{\theta}^{2}$ following from 98. The flow equation is given by
@@
in which the range of integration is $\Omega=\left[-\sqrt{k^{2}+\tilde{\theta}^{2}},\sqrt{k^{2}+\tilde{\theta}^{2}}\right]$. Performing the momentum integral and taking the real part, we obtain
@@
where $c^{(1)}_{k}$ and $c^{(2)}_{k}$ are defined in 86.
Now we solve the flow equation for $V_{k}(r^{2})$ with the UV boundary condition 94 at $k=\Lambda$. In the right-hand side panel of Figure 13, we show the result of the vacuum energy in terms of the $\theta$-parameter by setting $m=1$. We have not taken $r^{2}$-derivatives and hence the numerical solution cannot be extended easily beyond $\omega=2\pi$. As already discussed in D.1, this can be remedied by taking $r^{2}$-derivatives of 101. This goes beyond the scope of the present work.
The solution of the flow starts to deviate from the exact results with increasing $\theta$. However, the complex structure (’Silver Blaze’) of the theory is violated by the regulator and this may be the (only) reason for the deviation. Accordingly, this setup deserves further analysis using regulators that accommodate the complex structure. A more comprehensive analysis will be presented elsewhere.
Appendix E RG evolution of the effective potential
In Figure 14, we show the RG-evolution of the effective potential for different RG scales $k$. The cusps are given by the boundary lines of the different colored areas. The effective potential at any scale $k$ automatically takes the minimum value among every energy levels $n$ in the $\beta\rightarrow\infty$ limit, and thus the level-crossing lines would move as long as the flow down to IR scale. In both cases, we see a good agreement with the fRG-results with the results from the Schrödinger equation at small-$\theta$ regime, with an increasing but still small deviation for $\theta\gtrsim 9\pi$. This deviation originates in the simple LPA approximation. In Figure 15 we show the effective potential energy evaluated at $\theta=2\pi$ at different energy scales.
@@ Appendix F Flow of the cusp position
In this Appendix we derive relations for the flowing cusp positions. The derivation is reminiscent of that of the flow of the flowing solution of the equations of motion. We keep this derivations minimal as we do not use these relations in the present work. Still we have checked their applicability numerically.
@@
The numerator is simply the third derivative of the flow, evaluated at $\varphi_{k}^{\textrm{cusp}}$ with 102. All terms except one drop out because they are proportional to powers of $V_{k}^{(3)}(\varphi_{k}^{\textrm{cusp}})$.
@@
We emphasise that 104 holds true for any field value but for all but $\varphi^{\textrm{cusp}}$ both side in 105 vanish identically. In turn, for $\varphi_{k}^{\textrm{cusp}}$ they both are non-vanishing and we arrive at
@@
One may be tempted to use 106 in the finite-volume case where we have have $V^{+}_{k}=V^{-}_{k}$, but this leads us to 105 with vanishing left and right hand side, and 106 does not follow. Indeed, ignoring this fact leads us from 106 to
@@
Equation 107 vanishes on the cusps in clear contradiction to the observed flow of $\varphi_{k}^{\textrm{cusp}}$ either given by 103 or being extracted from the results. The vanishing of the right hand side of 107 originates in the fact that the flow of $V^{(1)}(\varphi_{k}^{\textrm{cusp}})$ is proportional to $V^{(3)}(\varphi_{k}^{\textrm{cusp}})=0$ and 107 vanishes. This concludes our discussion of the flow of $\varphi_{k}^{\textrm{cusp}}$.